OpenStudy (anonymous):

Y=Sin1/2x I need to graph it and give the amp and period. I don't understand it. I know it should be longer than y=sin2x but im not sure of the radians. Help?

OpenStudy (anonymous):

I think the amp is one and the period might be 2pi.

OpenStudy (anonymous):

@SolomonZelman I see you working on a sec/cos problem. Can you help me out?

OpenStudy (anonymous):

The 1/2 stretches the period from 2pi to 4pi and the amplitude stays the same. If the equation was y=2sin(x) then the amplitude would be 2.

OpenStudy (anonymous):

Really? Thats what i got.. but on the unit circle there is no 4pi.. so i thought that might be wrong..?

OpenStudy (anonymous):

You just go around the unit circle a second time.

OpenStudy (anonymous):

oh. Lol. Thanks! This was the only one i plugged into Wolfram Alpha that it couldn't understand and i didn't either. Wolfram just showed a straight line with weird values, lol.

OpenStudy (anonymous):



OpenStudy (anonymous):

yea thats what i got. And its amp is 1 and a period of 2pi right?

OpenStudy (anonymous):

When dealing with amplitude and period, you want to work with an x,y plane instead of the unit circle. The period starts, in this case at zero and ends where it begins. In the graph 2pi is only half way to the end of the period. 4pi is where it ends (and starts again).

OpenStudy (anonymous):

Since the amplitude wasn't modified, by multiplication of the sin, it stays at 1.

OpenStudy (anonymous):

mmm.. So the period is 4pi? Cuase my teacher said something like the graph repeates itself after 2pi again.

OpenStudy (anonymous):

y=sin(x) has a period of 2pi y=sin(2x) has a period of pi y=sin(1/2x) has a period of 4pi

OpenStudy (anonymous):

Oh. So question.. am I right to think of y=sin2x.. by multiplying it by 1/2? I feel like i know it doesn't add up but 1/2y=sinx

OpenStudy (anonymous):

Like, how am i supposed to know sin1/2x has a period of 4pi?

OpenStudy (anonymous):

the period is calculated as 2pi/B and the amplitude is A when graphing y=Asin(Bx) Since 2pi/(1/2) = 2pi x 2, then the period is 4pi

OpenStudy (anonymous):

Your equation was y=(1)sin(1/2x) where the 1 is assumed. A=1 B = 1/2 2pi/(1/2) = 4pi Gotta run to class myself, good luck!

OpenStudy (anonymous):

oh true. Thanks again! Thats actually sort of what i thought. lol. Yea gota go to class also.